# Projecting Gaussian Ellipsoids While Avoiding Affine Projection Approximation

### v1

@@ 4.2 Pre-filtering Strategy
The equation of the ellipsoid is $\mathcal{E}(\mathbf{x})=0$, where $\mathcal{E}(\mathbf{x})=(\mathbf{x}-\mathbf{p}_{c})^{T}\mathbf{\Sigma}_{c}^{-1}(\mathbf{x}-\mathbf{p}_{c})-9$. For the second type of Gaussian ellipsoid, we first need to find the lowest point of the ellipsoid surface in the direction of the z-axis, where the gradient direction of the ellipsoid surface is perpendicular to the plane $z=0$, satisfying $\nabla\mathcal{E}(\mathbf{x})=k\mathbf{n}$, in witch $\mathbf{n}=[0,0,1]$. By combining the ellipsoid equation and the gradient equation
@@ 5.1 Datasets and Implementation
Datasets  For training and testing, we perform experiments on images from a total of 16 real-world scenes. Specifically, we evaluate our ellipsoid-based projection method on 7 scenes from the Mip-NeRF360 dataset [3], 7 scenes from the Tanks&Temples dataset [25], and 2 scenes from the Deep Blending dataset [19]. The selected scenes showcase diverse styles, including bounded indoor environments and unbounded outdoor ones. To divide the datasets into training and testing sets, we follow the method used in 3DGS [24], assigning every 8th photo to the test set. For the images in the Mip-NeRF360 and Tanks&Temples datasets, we use $\frac{1}{2}$ original resolution for training and rendering.
Implementation  Our method is built on the open-source code of 3DGS [24]. Following the 3DGS framework, we set the number of training iterations to 30k for all scenes, using the same loss function and densification strategy as 3DGS, with all hyperparameters remaining consistent. We only modified the CUDA kernels in the projection parts of the forward and backward processes, replacing the Jacobian of the affine approximation of the projection transformation in 3DGS with our ellipsoid-based projection method. In the comparison experiments with Mip-Splatting [52], we integrated our ellipsoid-based projection method into Mip-Splatting in the same way. All our experiments are conducted on a single GTX RTX3090 GPU.
@@ 5.3 Comparisons with Mip-Splatting
To further demonstrate the effectiveness of our algorithm, we applied our ellipsoid-based projection method to Mip-Splatting [52] and compared it with the original Mip-Splatting method using PSNR, SSIM, and LPIPS metrics. We utilized Mip-Splatting’s core methods, 3D-filter and 2D-filter, but retained the Gaussian densification strategy consistent with 3DGS rather than adopting the newly proposed strategy from Mip-Splatting. In our experiments, we observed that the Gaussian densification strategy in Mip-Splatting led to a significant increase in the number of Gaussians for the original method, but had no noticeable effect when used with our projection transformation. Therefore, we conclude that this strategy is specific to Mip-Splatting. We set the 2D-filter parameter to 0.25, making the kernel size a $3\times 3$ pixel region on the image. As shown in the results in Tab. 3, applying the ellipsoid-based projection method to Mip-Splatting improved rendering quality. Similar to the results in Sec. 5.2, our method achieved overall metric improvements on the Mip-NeRF360 dataset, showed a slight decrease in PSNR on the Tanks&Temples dataset, and a slight decrease in SSIM on the Deep Blending dataset. Visual results in Fig. 4 show that, similar to 3DGS [24] in Sec. 5.2, Mip-Splatting with the ellipsoid-based projection method also reduces blurriness and artifacts, while enhancing scene consistency and details.
@@ 5.4 Limitations
Our method does not surpass 3DGS [24] in every metric across all scenes, indicating there is still room for improvement. In Sec. 5.3, we observed that Mip-Splatting’s [52] Gaussian densification strategy had limited impact on our method, suggesting that we may not be using the most suitable Gaussian densification strategy. If a more appropriate strategy could be identified, rendering quality might be further enhanced. In the experiments in Sec. 5.2 and Sec. 5.3, we used fixed values for the filter parameters. However, we found that our method is more sensitive to filter parameters than 3DGS, whether for the screen-space dilation filter in 3DGS or the 2D mip filter in Mip-Splatting. Additionally, the optimal filter parameters vary across different scenes, making parameter selection crucial. Since filter parameters are continuous values, it is difficult to identify optimal values through repeated experiments. One potential solution is to include the filter parameters as optimizable variables and optimize them during training.
@@ 6 Conclusion
We propose an ellipsoid-based projection method that avoids the errors introduced by the affine approximation of the projection transformation used in 3DGS. Additionally, we introduce a pre-filtering strategy to remove Gaussian ellipsoids that negatively impact the system or are outside the cone of vision, enhancing system robustness. Comparative experiments with 3DGS demonstrate that our method significantly improves rendering quality in both complex indoor and outdoor scenes, while also further accelerating rendering speed. By integrating our method with Mip-Splatting, rendering quality was further improved, proving the versatility of our method and its ease of adaptation to any work based on 3DGS.

### v3

@@ 4.2 Pre-filtering Strategy
The equation of the ellipsoid is $\mathcal{E}(\mathbf{x})=0$, where $\mathcal{E}(\mathbf{x})=(\mathbf{x}-\mathbf{p}_{c})^{T}\mathbf{\Sigma}_{c}^{-1}(\mathbf{x}-\mathbf{p}_{c})-9$. For the second type of Gaussian ellipsoid, we first need to find the lowest point of the ellipsoid surface in the direction of the z-axis, where the gradient direction of the ellipsoid surface is perpendicular to the plane $z=0$, satisfying $\nabla\mathcal{E}(\mathbf{x})=k\mathbf{n}$, where $\mathbf{n}=[0,0,1]$. By combining the ellipsoid equation and the gradient equation
@@ 5.1 Datasets and Implementation
Datasets  For training and testing, we perform experiments on images from a total of 16 real-world scenes. Specifically, we evaluate our ellipsoid-based projection method on 7 scenes from Mip-NeRF360 dataset [3], 7 scenes from Tanks&Temples dataset [25], and 2 scenes from Deep Blending dataset [19]. The selected scenes showcase diverse styles, including bounded indoor environments and unbounded outdoor ones. To divide the datasets into training and testing sets, we follow the method used in 3DGS [24], assigning every 8th photo to the test set. For images in Mip-NeRF360 and Tanks&Temples datasets, we use $\frac{1}{2}$ original resolution for training and rendering.
Implementation  Our method is built on the open-source code of 3DGS [24]. Following the 3DGS framework, we set the number of training iterations to 30k for all scenes, using the same loss function and densification strategy as 3DGS, with all hyperparameters remaining consistent. We only modified the CUDA kernels in the projection parts of the forward and backward processes, replacing the Jacobian of the affine approximation of the projection transformation in 3DGS with our ellipsoid-based projection method. In the comparison experiments with Mip-Splatting [52], we integrated the ellipsoid-based projection method into Mip-Splatting in the same way. All our experiments are conducted on a single GTX RTX3090 GPU.
@@ 5.3 Comparisons with Mip-Splatting
To further demonstrate the effectiveness of our algorithm, we applied our ellipsoid-based projection method to Mip-Splatting [52] and compared it with the original Mip-Splatting method using PSNR, SSIM, and LPIPS metrics. We utilized Mip-Splatting’s core methods, 3D-filter and 2D-filter, but retained the Gaussian densification strategy consistent with 3DGS rather than adopting the newly proposed strategy from Mip-Splatting. In our experiments, we observed that the Gaussian densification strategy in Mip-Splatting led to a significant increase in the number of Gaussians for the original method, but had no noticeable effect when used with our projection transformation. Therefore, we conclude that this strategy is specific to Mip-Splatting. We set the 2D-filter parameter to 0.25, making the kernel size a $3\times 3$ pixel region on the image. As shown in the results in Tab. 3, applying the ellipsoid-based projection method to Mip-Splatting improved rendering quality. Similar to the results in Sec. 5.2, our method achieved overall metric improvements on the Mip-NeRF360 dataset, showed a slight decrease in PSNR on the Tanks&Temples dataset, and a slight decrease in SSIM on the Deep Blending dataset. Visual results in Fig. 4 show that, similar to 3DGS [24] in Sec. 5.2, Mip-Splatting with the ellipsoid-based projection method also reduces blur and artifacts, while enhancing scene consistency and details.
@@ 5.4 Limitations
Our method does not surpass 3DGS [24] in every metric across all scenes, indicating there is still room for further improvement. In Sec. 5.3, we observed that Mip-Splatting’s [52] Gaussian densification strategy had limited impact on our method, suggesting that we may not be using the most suitable Gaussian densification strategy. If a more appropriate strategy could be identified, rendering quality might be further enhanced. In the experiments in Sec. 5.2 and Sec. 5.3, we used fixed values for the filter parameters. However, we found that our method is more sensitive to filter parameters than 3DGS, whether for the screen-space dilation filter in 3DGS or the 2D mip filter in Mip-Splatting. Additionally, the optimal filter parameters vary across different scenes, making parameter selection crucial. Since filter parameters are continuous values, it is difficult to identify optimal values through repeated experiments. One potential solution is to include the filter parameters as optimizable variables and optimize them during training.
@@ 6 Conclusion
We propose an ellipsoid-based projection method that avoids errors introduced by the affine approximation of the projection transformation used in 3DGS. Additionally, we introduce a pre-filtering strategy to remove Gaussian ellipsoids that negatively impact the system or are outside the cone of vision, enhancing system robustness. Comparative experiments with 3DGS demonstrate that our method significantly improves rendering quality in both complex indoor and outdoor scenes, and accelerates rendering speed at the same time. By integrating our method with Mip-Splatting, rendering quality was further improved, proving the versatility of our method and its ease of adaptation to any work based on 3DGS.